# Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory

### v1

@@ 3.2 Deep Pre-trained Time Series Features
Overview of Presto deep features. In contrast to harmonic features that fit mathematical curves to individual bands, we evaluate the effectiveness of deep features obtained by the pre-trained time series foundation model, which is Lightweight, Pre-trained Transformers for Remote Sensing Timeseries (Presto) developed and trained by Tseng et al. (2024) for general-purpose remote sensing time series classification and regression. This model is based on transformer architecture and was originally pretrained on 21.5 million pixel time-series with 12-month contiguous intervals and static topographic and coordinates data. Each month’s composite satellite data included S1, S2 and its NDVI, ERA5, Dynamic World (DW), extracted between 2020-01-01 and 2021-12-31. The static SRTM and coordinates data are added to the time-series data as input of the Presto encoder. The model was originally pretrained through Masked Auto Encoding (MAE) (He et al., 2022) reconstruct masked sequences across different satellite modalities. The advantages of this model include the ability to handle freely available multi-source and multi-temporal data even when some sources (e.g., ERA5) or temporal data (e.g., S2 values in November) are missing, and computational efficiency when processing large areas. We obtained 128 dimensional deep features from the encoder output.
@@
Multi Layer Perceptron (MLP) classifier was attached for fine-tuning of Presto encoder to the downstream task of tree species classification, which is described in section˜4. We used the fine-tuned encoder to extract deep features from the test set for the comparison between Harmonic and Seasonal Features and Presto deep features.
As shown in fig.˜2, we directly input 12-monthly time series ($T=12$ timesteps) for (i) 2 Sentinel-1 bands (S1), (ii) 10 Sentinel-2 bands (S2) and (iii) NDVI (NDVI), (iv) 2 ERA5 features (ERA5), (v) Dynamic world tensor with all tree classes value (equal to value of 1. DW). To every pixel-time series we appended two static-in-time products: (i) elevation and slope from the SRTM digital elevation model (Farr et al., 2007) (TG) and (ii) location coordinates (Loc) of each pixel as input tokens to fine-tuned Presto encoder. Location coordinates (Loc) is 3D static in time Cartesian coordinates computed from the latitude and longitude of the pixel’s geographical location.
@@
    $\mathbf{e}_{\text{S1}}$: S1 VV and VH bands.
@@
    $\mathbf{e}_{\text{S2-RGB}}$: S2 RGB(B2, B3, B4) bands.
@@
    $\mathbf{e}_{\text{S2-RE}}$: S2 Red Edge (B5, B6, B7) bands.
@@
    $\mathbf{e}_{\text{S2-NIR10}}$: S2 NIR (B8) 10m band.
@@
    $\mathbf{e}_{\text{S2-NIR20}}$: S2 NIR (B8A) 20m band.
@@
    $\mathbf{e}_{\text{S2-SWIR}}$: S2 SWIR(B11, B12) bands.
@@
    $\mathbf{e}_{\text{NDVI}}$: Computed from S2 B4 and B8.
@@
    Channel Group Encoding: A learnable encoding was added for each channel group $c\in\mathcal{C}=\{\text{S1},\text{S2 RGB},\ldots,\text{ERA5},\text{TG},\text{Loc}\}$, to capture source-specific information.
@@
Figure 2: Presto encoder feature extraction for downstream tasks. Figure modified from (Tseng et al., 2024).
@@
Data Normalization (for deep features). The data normalization follows the original Presto setup (Tseng et al., 2024), with adjustments based on the observed value ranges for each source. Sentinel-1 VH and VV bands, originally ranging from –31 to 17 dB, are normalized by shifting +25 and dividing by 25, resulting in a range of –0.24 to 1.68. Sentinel-2 spectral bands, with raw values from 10 to 15,769, are scaled by dividing by 10,000, yielding normalized values between 0.00 and 1.58. ERA5 temperature data, originally 278 to 295 K (equivalent to 6 to 23°C), are shifted by –272.15 and divided by 35, producing a range of 0.17 to 0.66. ERA5 precipitation, ranging from 0.008 to 0.208 m, is normalized by dividing by 0.03, resulting in values from 0.27 to 6.93. SRTM elevation data, spanning –27 to 331 m, are divided by 2,000 to give a range of –0.01 to 0.17, while slope values from 0.0 to 39.3° are scaled by dividing by 50, yielding normalized values between 0.00 and 0.79.
4 Experimental Setup
4.1 Comparative evaluation framework
  1. 1.
    Sensor contribution analysis: S1 vs S2 vs S1+S2 using harmonic features with RF
  2. 2.
    Feature representation comparison: Harmonic features vs deep features using RF
  3. 3.
    Classifier architecture evaluation: RF vs MLP using deep features
4.2 Model fine-tuning
For fine-tuning the pre-trained models, the encoder block from the pre-trained model was reused and trained with newly attached Multi Layer Perceptron (MLP) classifier for the downstream task. This fine-tuning model is trained on the 1,479 and 13,790 samples for 7 aggregated groups classification task, and the 1,462 samples for the 13 dominant species classification task. Pre-trained encoder was set as trainable in this process.
Based on the current state-of-the-art classifier of MLP on tree species classification (Mouret et al., 2024), this classifier architecture consists of a 3-layer MLP with no dropout (rate: 0.0). The layer configuration comprises 1024 nodes in the first layer, 512 nodes in the second layer, and 256 nodes for the last layer, with batch normalization and ReLU activation applied throughout. The training parameters are configured as follows: a learning rate of 0.0001 is used with a weight decay of 0.00746. The model is trained for a maximum of 100 epochs using a batch size of 64. The training process employs cross-entropy loss for optimization. The best fine-tuned models and its accuracy were selected based on best validation loss.
4.3 Data splitting strategy
Train-test split ratio is a critical factor in model performance on downstream tasks. If dataset size is large, 80/20 or higher ratios are commonly used (Rácz et al., 2021; Joseph, 2022). Based on our imbalanced dataset and previous research (Blickensdörfer et al., 2024), a 70/30 (approximately 2:1 ratio) train-test split was applied for fine-tuning and RF classification to increase model performance and reduce standard deviation on tree species classification over large areas for NFI. During fine-tuning and subsequent RF classification for features extracted from the fine-tuned model, the same training and test data were used. This practice ensures that extracted features are not trained on test data, which would cause data leakage. By adopting a class-per-split strategy, each class is equally split into train and test data at this ratio.
In remote sensing machine learning, ensuring that the train-test split strategy does not introduce spatial autocorrelation is crucial for obtaining unbiased model performance estimates Karasiak et al. (2022). Our NFI data accommodates a sampling strategy which selects each plot with a density of 1 point per 100ha (Schelhaas et al., 2014). This means that each random point is designated per square kilometer. Additionally, the Francini data removed adjacent pixels with a minimum distance of 15m (Hermosilla et al., 2022; Francini et al., 2024). These sampling strategies ensure that the train-test split does not introduce strong spatial autocorrelation in the data; therefore, no additional train-test split strategy was applied in this study.
4.4 Performance evaluation
Model performance was evaluated using a comprehensive set of established accuracy metrics: per-class precision and recall, F1 scores, confusion matrices, and overall accuracy. While overall accuracy provides a general assessment of classification performance, it exhibits reduced sensitivity to class imbalance issues that are prevalent in ecological datasets. Therefore, the F1 score, which represents the harmonic mean of precision and recall, was employed as a supplementary metric to provide a more balanced evaluation of model performance when addressing the inherent class imbalances in tree species distribution data. Confusion matrices were used to visualize the model’s performance at the class level. These metrics are widely used in tree species classification studies (Goutte and Gaussier, ; Hermosilla et al., 2022; Francini et al., 2024; Blickensdörfer et al., 2024) and provide comprehensive assessment of classification performance across all species classes.
4.5 Statistical validation
To evaluate variability of model performance, each model was evaluated five times with different random seeds, which affect to train-test split and model initialization. The mean and standard deviation of the accuracy metrics were calculated for each model. This approach provides a more robust evaluation of model performance.
5 Results
5.1 Influence of multi-source input features
Following the approach of Francini et al. (2024), we first evaluated the contribution of Sentinel-1 radar data to optical Sentinel-2 classification. Using RF with harmonic features, the addition of Sentinel-1 provided modest improvements: 0.4% for the 13-class COMB dataset and 1.4% for the 7-class SIMB dataset. These results justify the inclusion of Sentinel-1 features in our subsequent comparisons.
We first evaluated whether combining optical and radar satellite data improves tree species classification using hand-crafted features in a RF classifier. Feature sets were derived from seasonal medoid composites and harmonic fits, using either Sentinel-2 (S2) alone or in combination with Sentinel-1 (S1). The goal was to assess whether radar backscatter, which captures structural and moisture-related canopy properties, offers complementary information to optical reflectance data. As shown in Table 2 the inclusion of S1 provided modest but consistent improvements across both datasets. For the 13-class dataset (COMB), accuracy increased from 66.34% to 66.71% when adding S1 to S2-derived features. In the 7-class dataset (SIMB), accuracy improved from 72.87% to 74.22%. These results suggest that the radar signal contributes marginally to distinguishing between species, especially under a more simplified class grouping. While the absolute gains remain small, incorporating S1 ensures a more robust feature representation, particularly in settings where optical signals may be limited by cloud cover or seasonal variability. The modest accuracy gains observed here are also useful to contextualize the more substantial differences in performance reported in the next sections, which stem from the feature representations rather than the data sources themselves.
Label Features Sentinel-1 Sentinel-2 S1-S2 Combined
Band Acc. Band Acc. Band Acc.
COMB (13 classes) Seasonal (S) S1 51.03 S2 64.03 S1-S2 64.52
Harmonic (H) S1 55.04 S2 63.79 S1-S2 64.64
All (S+H) S1 57.72 S2 66.34 S1-S2 66.71
SIMB (7 classes) Seasonal (S) S1 55.60 S2 70.18 S1-S2 71.93
Harmonic (H) S1 60.19 S2 71.39 S1-S2 71.52
All (S+H) S1 62.75 S2 72.87 S1-S2 74.22
5.2 Deep vs Harmonic features
Table 3 demonstrates the superior performance of deep features over hand-crafted features when both are evaluated using RF classification. Deep features consistently outperform traditional harmonic and seasonal features across all three datasets, with F1 score improvements of 6-8 percentage points for the imbalanced datasets (COMB and SIMB) and over 10 percentage points for the balanced dataset (SIBA).
@@ 5.2 Deep vs Harmonic features
Table 3: Classification performance comparison (in %) across feature types (deep vs. harmonic+medoid) and classifiers (MLP vs. RF). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance among MLP and RF per dataset type.
Classification Results
Dataset type Data size MLP Classifier
(Deep features)
RF Classifier
(Deep features)
RF Classifier
(Harm.+med. Francini et al. (2024))
Overall Accuracy SIMB (7 classes) 1479 77.13 ± 1.17 74.62 ± 1.93 70.58 ± 1.78
COMB (13 classes) 1462 67.99 ± 1.33 66.77 ± 1.51 62.71 ± 1.51
SIBA (7 classes) 13,790 98.21 ± 0.20 95.28 ± 0.60 84.31 ± 0.99
F1 Score SIMB (7 classes) 1479 68.24 ± 2.85 60.54 ± 3.08 51.57 ± 3.36
COMB (13 classes) 1462 52.45 ± 2.25 46.84 ± 2.65 38.54 ± 2.14
SIBA (7 classes) 13,790 98.20 ± 0.20 95.27 ± 0.60 84.26 ± 0.99
For direct comparison using identical S1+S2 sensor inputs (detailed in Appendix Table A.1), deep features show F1 score improvements of 12.3 percentage points (SIMB), 9.7 percentage points (COMB), and 10.7 percentage points (SIBA), confirming consistent performance advantages across different input configurations.
We compared classification performance between traditional hand-crafted features and deep features extracted from a pre-trained transformer model. In both cases, a RF classifier was used to isolate the effect of the feature representation. As shown in Table 3, deep features extracted from the pre-trained Presto model consistently outperformed the harmonic and medoid predictors across all dataset.
On the 7-class dataset (SIMB), models trained with deep features achieved an overall accuracy of 74.62% (±1.93), compared to 70.58% (±1.78) for the harmonic+medoid approach – an improvement of approximately 4 percentage points. The difference was even more pronounced when examining F1 scores, where deep features yielded 60.54% (±3.08) versus 51.57% (±3.36) for traditional features, representing an improvement of nearly 9 percentage points. In the more complex 13-classes case (COMB), deep features kept the advantage over traditional hand-crafted features, achieving an overall accuracy of 66.77% (±1.51) compared to 62.71% (±1.51). Similarly, the F1 score showed a substantial improvement from 38.54% (±2.14) with traditional features to 46.84% (±2.65) with deep features – a gain of more than 8 percentage points. This indicates that the deep learning approach handles the increased complexity of multi-class classification more effectively. The highest performance gains were observed in the Francini dataset (SIBA), where deep features achieved a remarkable overall accuracy of 95.28% (±0.60), representing an improvement of nearly 11 percentage points over the traditional features’ accuracy of 84.31% (±0.99). The F1 scores demonstrated comparable improvements, with deep features achieving 95.27% (±0.60) compared to 84.26% (±0.99) for traditional features.
These results suggest that deep features extracted from the pre-trained transformer model are more effective at capturing species-specific spectral and temporal signatures than traditional hand-crafted features. The improvement is most pronounced in the SIBA dataset – where class separation is clearer due to the balanced classes and a higher sample size per class – but performance gains are evident across all dataset configurations examined in this study.
To evaluate the influence of classifier architecture, we compared the Multi Layer Perceptron (MLP) classifier for tree species classification Mouret et al. (2024) with a RF, both trained on the same deep features extracted from the pre-trained Presto model. This isolates the effect of the classifier while keeping the input representation fixed.
Across all datasets, the MLP classifier consistently outperformed the RF classifier in both overall accuracy and F1 score (Table 3). In the the 7-class dataset (SIMB), the MLP achieved an overall accuracy of 77.13% (±1.17) compared to 74.62% (±1.93) for the RF with a difference of approximately 2.5 percentage points. The performance gap was even more pronounced in terms of F1 score, with the MLP achieving 68.24% (±2.85) versus 60.54% (±3.08) for RF, representing a substantial improvement of nearly 8 percentage points. For the more challenging 13-class dataset (COMB), the MLP classifier maintained its advantage with an overall accuracy of 67.99% (±1.33) compared to 66.77% (±1.51) for the RF classifier. Similarly, the F1 score showed the MLP’s superior performance with 52.45% (±2.25) versus 46.84% (±2.65) for RF. The largest performance gain was observed in the balanced dataset (SIBA), where MLP reached 98.21% accuracy and 98.20% F1, compared to 95.28% and 95.27% with RF.
Notably, the MLP classifier not only achieved higher mean performance values but also demonstrated more consistent results across repeated runs, as evidenced by the lower standard deviations in most cases. These results indicate that while both classifiers can effectively leverage deep features, the MLP classifier with pre-trained Presto offers superior performance for forest type classification tasks across datasets of varying complexity and size. The improved class separation suggests that MLPs better exploit species-specific spectral and temporal signatures embedded in the deep features—patterns that may be less accessible to tree-based models like RF, especially in more balanced or data-rich scenarios.
The S1+S2 only comparison (detailed in Appendix Table A.2) confirms these findings with F1 score improvements of 6.7 percentage points for SIMB (70.57% vs 63.83%), 6.7 percentage points for COMB (54.91% vs 48.24%), and 3.1 percentage points for SIBA (98.07% vs 94.94%), demonstrating consistent MLP advantages across different input configurations.
5.3 Class level performances
To better understand class-level behavior, we visualized the confusion matrices of the best-performing model – an MLP fine-tuned on deep features – for all three datasets (COMB, SIMB, SIBA). These matrices are shown in Figure 3, with each cell reporting the number of predicted samples and the corresponding precision (i.e., percentage of correct predictions among all predicted samples for that class). In the 7-class NFI dataset (SIMB), the model performs strongly across most classes. Pinus shows high precision with 1,668 correctly predicted samples and a precision of 84.7%. Dark Conifer and Populus also achieve high precision with 1,448 samples at 73.5% and 1,833 samples at 93.0%, respectively. Broadleaved classes exhibit more overlap: Quercus (oak) has 1,478 samples with 75.0% precision, while Other Broadleaves and Fagus reach 71.6% and 91.6%, respectively. Confusion is mostly observed between Quercus, Populus, and Other Broadleaves, reflecting the spectral similarity within this group.
In the more complex 13-class dataset (COMB), misclassifications are more pronounced, but key species maintain strong separation. Pinus sylvestris achieves 77.4% precision, Pseudotsuga menziesii reaches 83.3%, and Picea spp. also stands at 83.3%. Populus spp. improves to 73.9%, and Betula spp. achieves 100% precision, though on a smaller sample size. In contrast, Quercus robur/petraea reaches only 53.9%, and Other Broadleaves suffers from overlapping predictions with Quercus and Populus. Despite the increased class count, precision remains relatively high for most coniferous classes and for broadleaves with distinct phenology.
In the balanced dataset (SIBA), the model achieves near-perfect class separation. All classes have 1,970 samples, and precision ranges between 84.0% (Larix) and 93.0% (Populus). Other classes include Fagus (91.6%), Pines (84.7%), Oak (75.0%), and Other Broadleaves (71.6%). The confusion is minimal and symmetrically distributed, confirming that the model benefits substantially from balanced data and larger sample size.
Figure 3: Confusion matrices for the best performing models with deep features by MLP classifier. Simplified & Imbalanced (SIMB)(left), Complex & Imbalanced (COMB)(center), Simplified & Balanced (SIBA) (right)
6 Discussion
6.1 Multi-modal sensor integration
Our findings on the modest but consistent benefits of combining Sentinel-1 radar with Sentinel-2 optical data align well with recent European tree species mapping studies. Blickensdörfer et al. (2024) demonstrated that Sentinel-1 SAR data effectively fills potential data gaps in cloud-affected regions, particularly important for national-scale applications where clear-sky observation density varies substantially. Similarly, Liu et al. (2023) confirmed that Sentinel-1 provides supplementary structural information to Sentinel-2 optical data, contributing to increased reliability of large-scale tree species mapping. Our results showing 0.4-1.4% accuracy improvements when adding Sentinel-1 to Sentinel-2 features corroborate these findings and provide confidence in the robustness of multi-modal approaches for operational NFI applications.
6.2 Deep vs hand-crafted features: an emerging paradigm
The transition from hand-crafted to learned feature representations marks a significant shift in remote sensing methodology. While deep learning has proven transformative in crop type mapping and land cover classification (Jakubik et al., 2023; Szwarcman et al., 2025; Wang et al., 2023), its application to tree species classification at NFI scale represents a relatively recent development. Our results demonstrate consistent improvements in F1 scores of 6-12 percentage points across datasets of varying class balance and complexity, suggesting that transformer-derived features capture phenological and structural variability more effectively than engineered metrics. This performance gain supports the ongoing shift in remote sensing from task-specific feature engineering toward transferable, learned representations that can adapt to diverse ecological contexts without domain-specific tuning.
6.3 Neural network architectures for learned features
The systematic superiority of MLP classifiers over RF when both were trained on deep features reflects broader patterns observed in high-dimensional, learned feature spaces. Neural architectures tend to better exploit the latent structure present in deep representations, where complex non-linear relationships between features may not be effectively captured by tree-based models. This finding aligns with recent advances in representation learning, where the pairing of learned features with compatible downstream architectures has proven critical for optimal performance (Mouret et al., 2024).
6.4 Computational efficiency and operational scalability
6.5 Data quality and class balance considerations
Our comparative analysis across datasets of varying quality and balance reveals critical insights for operational NFI implementations. The stark performance differences between the NFI datasets (COMB and SIMB) and the Francini dataset (SIBA) underscore the fundamental importance of training data characteristics for deep learning approaches.
The Francini dataset’s superior performance (98.2% F1 score) compared to the imbalanced NFI datasets (52.5-68.2% F1 scores) primarily reflects the impact of class balance rather than dataset size alone. While the Francini dataset benefits from both larger sample size (13,790 vs 1,462-1,479) and better class balance, the magnitude of improvement suggests that addressing class imbalance represents a critical optimization opportunity for operational systems.
This finding has important implications for NFI data collection strategies. Resources may be more effectively allocated to targeted collection for underrepresented species classes. The approach demonstrated by Francini et al. (2024), which supplements ground-truth NFI data with additional samples through visual interpretation of high-resolution satellite imagery, offers a practical pathway for addressing these imbalances. Tools such as Collect Earth Online (Saah et al., 2019) provide accessible platforms for additional data collection.
6.6 Limitations and future directions
Several limitations merit consideration for operational deployment. First, deep features were extracted from a model pre-trained on a limited temporal and spatial range (2020-2021 global data), and further domain adaptation might improve performance for underrepresented classes to country’s forests. Second, while MLPs demonstrated superior performance in our controlled experiments, their generalization in operational settings with smaller or more imbalanced training sets requires further investigation.
Future research directions include: (i) adaptation to new regions or tree species not present in the training data, testing the transferability of learned representations across different forest ecosystems; (ii) integration of additional modalities such as hyperspectral or LiDAR data where available, potentially capturing structural metrics that complement spectral-temporal patterns; and (iii) development of computational efficient uncertainty quantification methods to assess classification confidence for operational decision-making.
7 Summary and conclusions
This study examined the efficacy of deep learning approaches for tree species classification in National Forest Inventories, yielding several significant findings. First, deep features extracted from pre-trained Presto models consistently outperformed traditional methods employing harmonic and medoid predictors across varying dataset sizes and classification complexities. This performance advantage was particularly evident with larger and more balanced training datasets, as demonstrated by the results from the Francini dataset, indicating superior scalability of deep learning approaches with increased high-quality data availability. Second, the MLP classifier demonstrated superior performance compared to RF in our experimental setting, corroborating previous findings (Mouret et al., 2024). Collectively, these results demonstrate the potential for implementing computationally efficient, large-scale tree species mapping within NFI systems.
This research significantly advances our understanding of deep learning applications in forestry by having demonstrated the effectiveness of fine-tuning pre-trained deep learning models for tree species classification, establishing a computationally efficient approach for improving NFI accuracy.
CRediT authorship contribution statement
Takayuki Ishikawa: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis, Conceptualization. Carmelo Bonannella: Writing – review & editing, Supervision, Conceptualization. Bas J. W. Lerink Writing – review & editing, Data curation. Marc Rußwurm Writing – review & editing, Supervision, Conceptualization.
Declaration of competing interest
Acknowledgement
This research is based on data from the Dutch National Forest Inventory (NFI), commissioned by the Ministry of Agriculture, Nature and Food Quality of the Netherlands. We thank the Ministry and the agencies responsible for collecting and providing these nationally representative forest data, which underpin our analysis. C.B. (Carmelo Bonannella) was supported by the Open-Earth-Monitor Cyberinfrastructure project, funded by the European Union’s Horizon Europe research and innovation programme [grant agreement No.101059548].
Data availability
The source code is available on:
Appendix A Appendix
A.1 Additional results
A.1.1 Results of models using only S1 and S2 data for fine-tuning
Our primary models incorporated all available satellite data as input for both the second stage of pre-training and fine-tuning to extract deep features. In this section, we present results from models using only S1 and S2 data as input for fine-tuning and feature extraction, excluding SRTM and ERA5 data.
Deep features model performance
Table A.1 demonstrates that deep features extracted by pre-trained Presto from only S1 and S2 data outperformed harmonic and medoid features in the RF classifier. These results were comparable to models that utilized all satellite data as input.
Table A.1: Classification performance comparison between deep features and harmonic+medoid features (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between deep features extracted from Presto and harmonic+medoid features per dataset type.
@@ A.1.1 Results of models using only S1 and S2 data for fine-tuning
Dataset type Data size Deep feat.
((Tseng et al., 2024))
@@
Overall Accuracy NFI data (7 classes) 1480 75.47 ± 1.15 70.58 ± 1.78
NFI data (13 classes) 1462 67.77 ± 2.38 62.71 ± 1.51
Francini data (7 classes) 13,790 94.96 ± 0.58 84.31 ± 0.99
F1 Score NFI data (7 classes) 1480 63.83 ± 1.21 51.57 ± 3.36
NFI data (13 classes) 1462 48.24 ± 4.31 38.54 ± 2.14
Francini data (7 classes) 13,790 94.94 ± 0.58 84.26 ± 0.99
Table A.2 shows that the MLP classifier consistently achieved higher accuracy than the RF classifier, even when using only S1 and S2 data as input for the pre-trained Presto model.

### v2

@@ A.1 Implementation Details on Presto Embeddings
Calculation of Presto deep embeddings. In contrast to harmonic features that fit mathematical curves to individual bands, we evaluate the effectiveness of deep embeddings obtained by the pre-trained time series foundation model, which is Lightweight, Pre-trained Transformers for Remote Sensing Timeseries (Presto) developed and trained by Tseng et al. (2024) for general-purpose remote sensing time series classification and regression. This model is based on transformer architecture and was originally pretrained on 21.5 million pixel time-series with 12-month contiguous intervals and static topographic and coordinates data. Each month’s composite satellite data included Sentinel-1 (S1), S-2 and its NDVI, European Centre for Medium-Range Weather Forecasts Re Analysis v5 (ERA5), Dynamic World (DW), extracted between 2020-01-01 and 2021-12-31. The static Shuttle Radar Topography Mission (SRTM) and coordinates data are added to the time-series data as input of the Presto encoder. The model was originally pretrained through Masked Auto Encoding (MAE) (He et al., 2022) reconstruct masked sequences across different satellite modalities. The advantages of this model include the ability to handle freely available multi-source and multi-temporal data even when some sources (e.g., ERA5) or temporal data (e.g., S-2 values in November) are missing, and computational efficiency when processing large areas. We obtained 128 dimensional deep embeddings from the encoder output.
@@
MLP classifier was attached for fine-tuning of Presto encoder to the downstream task of tree species classification, which is described in section˜4. We used the fine-tuned encoder to extract deep embeddings from the test set for the comparison between Harmonic and Seasonal Features and Presto deep embeddings.
We directly input 12-monthly time series ($T=12$ timesteps) for (i) 2 Sentinel-1 bands (S-1), (ii) 10 Sentinel-2 bands (S-2) and (iii) NDVI, (iv) 2 ERA5 features (ERA5), (v) Dynamic World land cover class probabilities (DW). To every pixel-time series we appended two static-in-time products: (i) elevation and slope from the SRTM digital elevation model (Farr et al., 2007) (TG) and (ii) location coordinates (Loc) of each pixel as input tokens to fine-tuned Presto encoder. Location coordinates (Loc) are 3D static in time Cartesian coordinates computed from the latitude and longitude of the pixel’s geographical location.
@@
    $\mathbf{e}_{\text{S1}}$: S-1 VV and VH bands.
@@
    $\mathbf{e}_{\text{S2-RGB}}$: S-2 RGB(B2, B3, B4) bands.
@@
    $\mathbf{e}_{\text{S2-RE}}$: S-2 Red Edge (B5, B6, B7) bands.
@@
    $\mathbf{e}_{\text{S2-NIR10}}$: S-2 NIR (B8) 10m band.
@@
    $\mathbf{e}_{\text{S2-NIR20}}$: S-2 NIR (B8A) 20m band.
@@
    $\mathbf{e}_{\text{S2-SWIR}}$: S-2 SWIR(B11, B12) bands.
@@
    $\mathbf{e}_{\text{NDVI}}$: Computed from S-2 B4 and B8.
@@
    Channel Group Encoding: A learnable encoding was added for each channel group $c\in\mathcal{C}=\{\text{S-1},\text{S-2 RGB},\ldots,\text{ERA5},\text{TG},\text{Loc}\}$, to capture source-specific information.
@@
Data Normalization (for deep embeddings). The data normalization follows the original Presto setup (Tseng et al., 2024), with adjustments based on the observed value ranges for each source. S-1 VH and VV bands, originally ranging from –31 to 17 dB, are normalized by shifting +25 and dividing by 25, resulting in a range of –0.24 to 1.68. Sentinel-2 spectral bands, with raw values from 10 to 15,769, are scaled by dividing by 10,000, yielding normalized values between 0.00 and 1.58. ERA5 temperature data, originally 278 to 295 K (equivalent to 6 to 23°C), are shifted by –272.15 and divided by 35, producing a range of 0.17 to 0.66. ERA5 precipitation, ranging from 0.008 to 0.208 m, is normalized by dividing by 0.03, resulting in values from 0.27 to 6.93. SRTM elevation data, spanning –27 to 331 m, are divided by 2,000 to give a range of –0.01 to 0.17, while slope values from 0.0 to 39.3° are scaled by dividing by 50, yielding normalized values between 0.00 and 0.79.
A.2 Details of methodology
A.2.1 Data Splitting and training pipeline
The detailed breakdown of class distribution across train, validation, and test splits is provided in table˜A.1. For fine-tuning Presto, we used MLP-Train data as training data for the MLP classifier and evaluated loss values based on MLP-validation data.
Table A.1: Species/Group class distribution across train, validation, and test splits for three datasets.
Dataset Class Train Test Total
(train total) MLP-Train MLP-Val
NFI (dominantSpecies) Alnus spp 21 16 5 9 30
Betula spp 40 32 8 18 58
Fagus spp 40 32 8 18 58
Fraxinus spp 28 22 6 12 40
Larix spp 39 31 8 17 56
Other Pinus 62 49 13 27 89
Other Quercus 23 18 5 10 33
Other broadleaved 71 56 15 31 102
Picea spp 46 36 10 20 66
Pinus sylvestris 359 287 72 154 513
Populus spp 50 40 10 22 72
Pseudotsuga menziesii 62 49 13 28 90
Quercus robur petraea 178 142 36 77 255
Total 1,019 810 209 443 1,462
NFI (Group_F) Beech 40 32 8 18 58
Dark Conifer 112 89 23 48 160
Larix 39 31 8 17 56
Other Broadleaves 169 135 34 73 242
Pinus 422 337 85 181 603
Populus 50 40 10 22 72
Quercus 201 160 41 87 288
Total 1,033 824 209 446 1,479
Francini (group2) Beech 1,379 1,103 276 591 1,970
Dark Conifer 1,379 1,103 276 591 1,970
Larix 1,379 1,103 276 591 1,970
Other Broadleaf 1,379 1,103 276 591 1,970
Pinus 1,379 1,103 276 591 1,970
Populus 1,379 1,103 276 591 1,970
Quercus 1,379 1,103 276 591 1,970
Total 9,653 7,721 1,932 4,137 13,790
A.3 Additional results
A.3.1 Results of models using only S-1 and S-2 data for fine-tuning and embeddings extraction
Our primary models incorporated all available satellite data as input for fine-tuning and extraction of deep embeddings from Presto. In this section, we present results from models using only S-1 and S-2 data as input for fine-tuning and embeddings extraction, excluding SRTM and ERA5 data.
deep embeddings model performance
TableA.2 reports classification performance in RF classifier when fine-tuning Presto using only S-1 and S-2 data for fine-tune and embeddings extraction, excluding auxiliary topographic (SRTM) and climatic (ERA5) inputs. Across all datasets, fine-tuned Presto embeddings substantially outperform harmonic + medoid features for all evaluation metrics, highlighting the strong representational capacity of deep embeddings even when limited to optical and radar data. These results indicate that S-1 + S-2 only fine-tuning Presto embeddings remains highly effective and robust.
@@ A.3.1 Results of models using only S-1 and S-2 data for fine-tuning and embeddings extraction
Table A.2: Classification performance comparison between fine-tuned Presto embeddings using only S-1 and S-2 inputs and harmonic + medoid features (in %). Results are reported for a RF classifier. Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between the models per dataset type.
@@
Dataset type Data size Fine-tune Presto
embeddings
@@
Overall Accuracy SIMB (7 classes) 1,479 74.01 ± 1.83 69.48 ± 0.73
COMB (13 classes) 1,462 65.42 ± 1.92 62.80 ± 0.91
SIBA (7 classes) 13,790 92.31 ± 0.18 84.47 ± 0.41
F1 Macro SIMB (7 classes) 1,479 60.00 ± 2.12 51.03 ± 1.71
COMB (13 classes) 1,462 44.67 ± 3.25 38.95 ± 2.02
SIBA (7 classes) 13,790 91.95 ± 0.24 84.40 ± 0.40
F1 Weighted SIMB (7 classes) 1,479 72.96 ± 1.87 66.98 ± 1.06
COMB (13 classes) 1,462 61.32 ± 2.10 57.13 ± 1.25
SIBA (7 classes) 13,790 92.23 ± 0.20 84.40 ± 0.40
Table A.3 compares the performance of MLP and RF classifiers using fine-tuned Presto embeddings when only S-1 and S-2 data are provided as input. The MLP classifier consistently outperforms RF on the SIMB and COMB datasets across all metrics, with particularly notable improvements in F1 macro and weighted scores for the more class-imbalanced COMB dataset. In contrast, performance differences between the two classifiers are marginal for the SIBA dataset, where both models achieve similarly high accuracy and F1 scores. This suggests that classifier choice has a stronger impact under reduced input modalities and class imbalance, while its effect diminishes for larger and more balanced datasets.